The heat of combustion of propanol (C3H9O) at 298.15 K, determined in a bomb

calorimeter is 43.5 kJ/g. For the combustion of one mole of propanol, determine ΔH

(kJ/mol)?


1
Expert's answer
2021-12-21T00:40:02-0500

C3H8O ( l ) + 9/2 O2 ( g ) = 3 CO2 ( g ) + 4 H2O ( l )


ng = 3 - 9 / 2 = - 3 / 2 .


H = U + ng . R .T


U =- 43.5 kj / mol. ( given ) .

U = - 43.5 kj / mol.

U = - 43500 j / mol .

H = - 43500 j /mol + ( -3/2 ) × 8.314 × 298.15

H = ( - 43500 - 3718.22 ) j / mol.

H = - 47218.22 j / mol .


Combustion is relese energy so - ve .


In bomb calorimetre intetnal energy calculated .